# FaceSORT: a Multi-Face Tracking Method based on Biometric and Appearance Features

### v1

@@ Abstract
Tracking multiple faces is a difficult problem, as there may be partially occluded or lateral faces. In multiple face tracking, association is typically based on (biometric) face features. However, the models used to extract these face features usually require frontal face images, which can limit the tracking performance. In this work, a multi-face tracking method inspired by StrongSort, FaceSORT, is proposed. To mitigate the problem of partially occluded or lateral faces, biometric face features are combined with visual appearance features (i.e., generated by a generic object classifier), with both features are extracted from the same face patch. A comprehensive experimental evaluation is performed, including a comparison of different face descriptors, an evaluation of different parameter settings, and the application of a different similarity metric. All experiments are conducted with a new multi-face tracking dataset and a subset of the ChokePoint dataset. The ‘Paris Lodron University Salzburg Faces in a Queue’ dataset consists of a total of seven fully annotated sequences (12730 frames) and is made publicly available as part of this work. Together with this dataset, annotations of 6 sequences from the ChokePoint dataset are also provided.
@@ 1 Introduction
In this work, we propose FaceSORT, a new TbD based multi-face tracker inspired by StrongSORT. Instead of utilizing features extracted from two modalities (e.g., face and body)[42], two different types of features are extracted from the same modality (i.e., detected face). By using two different features from the same image patch, the required presence of body and face, which hardly exists in the considered scenario, is eliminated. The two features (i.e., face features and appearance features) are combined into a single cost value, which forms the basis for association. To the best of our knowledge, FaceSORT is the first multi-face tracking method that combines two features extracted from the same image patch. All other described multi-face trackers [24, 25, 31, 32, 33, 34, 35] rely solely on face features. The combination of the two features helps to better handle lateral or partially occluded faces. FaceSORT is evaluated based on a new multi-face tracking dataset. This dataset reflects the assumed scenario and is released as part of this work. To the best of our knowledge, the published dataset is the first multi-face tracking dataset in which people move in an (unorganized) queue towards a camera (gate). There is only one similar dataset, the ChokePoint dataset [47], in which several people move through a portal. In the released dataset, however, a queue is simulated in which people are talking, eating, pushing, shoving, etc.
@@
  • A novel multi-face tracking method (FaceSORT) is proposed that combines two different features (i.e., biometric (face) features and appearance features) extracted from the same image patch.
@@
The remainder of this paper is organized as follows: the proposed FaceSORT is described in section 2. Section 3 provides an overview of the proposed dataset. The experimental evaluation is described in section 4. An ablation study is conducted in section 5 and the key insights are summarized in the last section 6.
2 FaceSORT
FaceSORT is a multi-face tracker based on the basic variant of StrongSORT [17]. Since common face recognition models are usually trained on constrained face images, successful association (of detected faces and active traces) of lateral or partially occluded faces based on extracted face features can be difficult. To mitigate this problem, FaceSORT combines two different types of features extracted from the same image patch (i.e., detected faces). These features are: (i) face features (representing biometric information), and (ii) appearance features (representing visual appearance). The combination of these features is performed on the cost matrix level.
In common TbD based methods, the association relies on a cost matrix $C^{n\times m}$, where $n$ is the number of active trajectories (tracks) and $m$ is the number of detected objects. An element $C_{i,j}$ represents the cost of matching the $i$-th active track to the $j$-th detected object. The main objective of association is to find bipartite matches where the global cost is minimum (i.e., usually solved by the Hungarian algorithm [14]). In principle, $C_{i,j}$ is based on a distance $d$ from a stored/predicted cue of the $i$-th active track to a calculated/observed cue of the $j$-th detected object and is calculated in FaceSORT as follows,
@@ 2 FaceSORT
Figure 1: Overview FaceSORT association.
In DeepSORT, a matching cascade is applied in which the matching is performed in sequence, starting with the tracks that were correctly matched in the previous frame and ending with the tracks that have not been matched for the longest time. The matching cascade reduces the search space and could be beneficial when long occlusions are allowed. An overview of the core association procedure of FaceSORT is depicted in Figure 1. For all unmatched detections (after the matching cascade), an IoU-based association (as proposed in SORT [15]) is performed as a fallback routine. The general steps that are performed in FaceSORT for each frame are summarized in Algorithm 1.
Algorithm 1 FaceSORT at frame $t$
@@ 2.1 Computational Complexity
As shown in Algorithm 1, FaceSORT essentially comprises the matching cascade, the IoU fallback matching, the feature updating and the position prediction. In general, the complexity of finding bipartite matches using the Hungarian algorithm is $O(|\Phi|^{2}\times|\Psi|)$. This is also the worst case with cascade matching. In the considered scenario, it is very likely that the detected face $\Psi$ in frame $t$ were also present in frame $t-1$. Thus, it is reasonable that the complexity can be reduced by the matching cascade (i.e., to $O((\nicefrac{{|\Phi|}}{{x}})^{2}\times\Psi)$). For all unmatched detections after the matching cascade, the complexity of the IoU fallback matching is in worst case again $O(|\Phi|^{2}\times|\Psi|)$. The complexity for updating the memory for biometric and appearance features ($\Phi_{bio}$ and $\Phi_{app}$) is $O(|\Psi|)$ in each case. Finally, predicting the positions in frame $t+1$ using the Kalman filter has basically a complexity of $O(\Psi)$. Thus, the overall complexity of FaceSORT is,
@@ 4 Experimental Evaluation
FaceSORT is evaluated based on the proposed PLUSFiaQ (see section 3) and the ChokePoint[47] dataset. In the ChokePoint dataset, people moving through a portal (chokepoint) are recorded. The dataset comprises sequences from 18 different runs (i.e., different portal, session or walking direction), with 3 different camera perspectives available for each run (i.e., $18\times 3=54$ available sequences). Of these 54 sequences, only 6 sequences show multiple persons per frame, which is similar to the assumed scenario in this paper. These 6 sequences are therefore used for evaluation. However, no annotations are available for these 6 sequences and are created accordingly. In the ChokePoint dataset, people are reflected in the open glass door as they walk through the portal. These refelections are annotated as non-target classes. The annotations created have the same format as described in section 3 and are made available along with the PLUSFiaQ dataset.
@@ 4.1 Tracker Implementation
FaceSORT is implemented based on the StrongSORT implementation [^4] https://github.com/dyhBUPT/StrongSORT. The parameter $\alpha$ for the EMA feature updating strategy is set to $\alpha=0.9$. The weighting parameter $\beta$ (for weighting the contribution of biometric and appearance similarity compared to spatial distance) is set to 0.98. This gives the biometric and appearance similarity the most weight. $\theta_{pos}$ is set to 9.4877 (the 0.95 quantile of the chi-squared distribution with 4 degrees of freedom). These parameters are set according to [17]. The general cost threshold $\theta$ is set to 0.2 and the minimum IoU (fallback routine) to 0.3. For all experiments, $N_{init}=1$ and the maximum age parameter is set to $N_{max}=100$ frames (i.e., all tracks that were not successfully matched for more than 100 frames are deleted). With this setting, most occlusions are taken into account while avoiding interference from inter-track occlusions (as shown in Figure 5). As face detector yolov8555https://github.com/ultralytics/ultralytics with a confidence score threshold of 0.4 is used.
@@ 4.4 Experimental Results
In FaceSORT, the selection of the weighting parameter $\lambda$ (combination of biometric and appearance features) and the association threshold $\theta$ is crucial. Thus, the parameter selection is evaluated in detail along with different face features and a different similarity metric (Euclidean distance instead of cosine similarity).
@@ 4.4.4 Matching Score Distribution
A multi-face tracker can be compared with a biometric system (e.g., a set of query images (detected faces) are matched against a database of enrolled identities (active tracks)). Matching is based on a similarity score, and in a perfect system, the genuine scores do not overlap with the impostor scores. In the context of FaceSORT, the genuine scores represent the cosine similarity that result when a detected face is compared with the corresponding stored faces (active tracks). The imposter scores, on the other hand, represent the cosine similarities that are achieved when a detected face is compared with all other stored faces (tracks). Table 4 shows the mean and the standard deviation of the genuine and the imposter score distribution as well as their overlap (IoU). Considering the IoU, the most distinct face descriptor (lowest IoU) is SFace. Other face descriptors with a low IoU are Facenet512, VGG-Face, ArcFaceII and Dlib. These face descriptors (except Dlib) are also the face descriptors achieving the highest AssPr (see Figure 7). However, based on the IoU results, it is surprising that such low AssPr values are obtained with the Dlib face descriptor. A reason for the low AssPr values could be that, with a mean value of 0.07, the imposter scores are well below the threshold $\theta=0.2$ (applied to reject uncertain (imposter) matches). Thus, a closer look is taken at the selection of the threshold value $\theta$.
@@ 4.4.7 Compare Tracker
Table 8: Compare FaceSORT with state-of-the-art multi-face tracker, where MC and IoU denote FaceSORT wihtout matching cascade (MC) and/or IoU fallback matching.
@@
FaceSORT MC 0.5940 0.7147 0.7025 0.0169 0.8274 0.8603 0.8700 0.0107 0.0405 0.1776 0.0875 0.0400
FaceSORT IoU 0.5764 0.7033 0.6760 0.0187 0.6933 0.7863 0.7671 0.0283 0.0363 0.1676 0.0830 0.0494
FaceSORT MC IoU 0.5589 0.6924 0.6666 0.0292 0.6697 0.7721 0.7521 0.0392 0.0359 0.1664 0.0812 0.0618
In table 8 FaceSORT is compared with different state-of-the-art multi-face trackers. A comparison with StrongSORT including MC (i.e., when $\lambda$ is equal 1.0 or 0.0) is already performed in ‘1) Selection of a fixed $\lambda$’. The FaceSORT results are obtained with SFace features, $\lambda=0.1$ and $\theta=0.2$. Implementations are available for the MCFDR [^7] https://github.com/yjwong1999/OpenVINO-Face-Tracking-using-YOLOv8-and-DeepSORT and UMA [^8] https://github.com/yinjunbo/UMA-MOT. However, no implementation is available for the OMTMCM multi-face tracker. For this reason, OMTMCM is carefully implemented according to the description in [42]. To verify the implementation, the results for the MusicVideo dataset [70] are also presented in table 8. The MusicVideo dataset consists of 8 music videos (i.e., Apink, BrunoMars, Darling, GirlsAloud, HelloBubble, PussycatDolls, Tara and Westlife) from YouTube, is publicly available [^9] https://sites.google.com/site/shunzhang876/eccv16-face-tracking and was used in [42] to evaluate the proposed OMTMCM multi-face tracker. Since music videos are unconstrained videos (with different scenes, moving camera, etc.), they do not correspond to the considered scenario in which people move towards a gate for which FaceSORT is designed. For this dataset, OMTMCM performed best. However, in the considered sceneraio where people are moving towards a gate (PLUSFiaQ and ChokePoint), FaceSORT clearly outperformed the other evaluated state-of-the-art trackers.
@@ 5 Ablation Study
The main components of FaceSORT are the combination of biometric and appearance features, the matching cascade and IoU fallback matching. Running FaceSORT with either biometric (i.e., $\lambda=1.0$) or appearance features (i.e., $\lambda=0.0$) has already been evaluated in ‘1) Selection of a fixed $\lambda$’ of subsection 4.4.
Table 9: Frames per second for FaceSORT (FS), FaceSORT without the matching cascade (FS MC) and OMTMCM [42].
@@
In the matching cascade (described in section 2), the tracks that were successfully matched in the previous frame are considered first. This can reduce the search space $\Phi$ and can be particularly beneficial if long occlusions are allowed (i.e., a large active track memory $\Phi$). For all experiments, occlusions up to 100 frames are allowed (i.e., $N_{max}=100$). In table 9, execution times in frames per second (fps) are reported for FaceSORT with and without the matching cascade (MC) and compared with the re-implementation of a state-of-the-art tracker (i.e., OMTMCM [42]). All times were recorded on a standard desktop PC (i.e., Intel Core i5 gen9) and without taking face detection and feature extraction into account. It can be seen that the fps increases slightly when MC is applied. This is also reflected by the observation that on average less than half of all stored active tracks are required to successfully match all detected faces. However, applying the MC carries the risk that not the best matches are found. For example, it could happen that the best match (lowest cost) for a detected face is with a track that could not be matched in the last frames (e.g., a person returning from an occlusion), but the detected face is matched with a track that is tried in the MC before. The achieved tracking performance scores for FaceSORT (with SFace features, $\lambda=0.1$ and $\theta=0.2$) without the matching cascade (MC) are reported in table 8. Compared to results when using the matching cascade (table 8 first row) a performance decrease can be observed.
IoU fallback matching is applied to all unmatched detections after the matching cascade, including matching with tentative tracks (only confirmed tracks are considered in the matching cascade). To achieve the reported tracking performance of FaceSORT (table 8 first row), IoU matching is performed for 3.91%, 4.9% and 8.42% of all detected faces of the PLUSFiaQ, ChokePoint and MusicVideo dataset respectively. When disabling IoU fallback matching (IoU) the performance decreases significantly (see table 8). The performance achieved when FaceSORT is applied without matching cascade and IoU fallback matching is reported in table 8 (last row).

### v2

# FaceQSORT: a Multi-Face Tracking Method based on Biometric and Appearance Features

@@ Abstract
Tracking multiple faces is a difficult problem, as there may be partially occluded or lateral faces. In multiple face tracking, association is typically based on (biometric) face features. However, the models used to extract these face features usually require frontal face images, which can limit the tracking performance. In this work, a multi-face tracking method inspired by StrongSort, FaceQSORT, is proposed. To mitigate the problem of partially occluded or lateral faces, biometric face features are combined with visual appearance features (i.e., generated by a generic object classifier), with both features are extracted from the same face patch. A comprehensive experimental evaluation is performed, including a comparison of different face descriptors, an evaluation of different parameter settings, and the application of a different similarity metric. All experiments are conducted with a new multi-face tracking dataset and a subset of the ChokePoint dataset. The ‘Paris Lodron University Salzburg Faces in a Queue’ dataset consists of a total of seven fully annotated sequences (12730 frames) and is made publicly available as part of this work. Together with this dataset, annotations of 6 sequences from the ChokePoint dataset are also provided.
@@ 1 Introduction
In this work, we propose FaceQSORT, a new TbD based multi-face tracker inspired by StrongSORT. Instead of utilizing features extracted from two modalities (e.g., face and body)[42], two different types of features are extracted from the same modality (i.e., detected face). By using two different features from the same image patch, the required presence of body and face, which hardly exists in the considered scenario, is eliminated. The two features (i.e., face features and appearance features) are combined into a single cost value, which forms the basis for association. To the best of our knowledge, FaceQSORT is the first multi-face tracking method that combines two features extracted from the same image patch. All other described multi-face trackers [24, 25, 31, 32, 33, 34, 35] rely solely on face features. The combination of the two features helps to better handle lateral or partially occluded faces. FaceQSORT is evaluated based on a new multi-face tracking dataset. This dataset reflects the assumed scenario and is released as part of this work. To the best of our knowledge, the published dataset is the first multi-face tracking dataset in which people move in an (unorganized) queue towards a camera (gate). There is only one similar dataset, the ChokePoint dataset [47], in which several people move through a portal. In the released dataset, however, a queue is simulated in which people are talking, eating, pushing, shoving, etc.
@@
  • A novel multi-face tracking method (FaceQSORT) is proposed that combines two different features (i.e., biometric (face) features and appearance features) extracted from the same image patch.
@@
The remainder of this paper is organized as follows: the proposed FaceQSORT is described in section 2. Section 3 provides an overview of the proposed dataset. The experimental evaluation is described in section 4. An ablation study is conducted in section 5 and the key insights are summarized in the last section 6.
2 FaceQSORT
FaceQSORT is a multi-face tracker based on the basic variant of StrongSORT [17]. Since common face recognition models are usually trained on constrained face images, successful association (of detected faces and active traces) of lateral or partially occluded faces based on extracted face features can be difficult. To mitigate this problem, FaceQSORT combines two different types of features extracted from the same image patch (i.e., detected faces). These features are: (i) face features (representing biometric information), and (ii) appearance features (representing visual appearance). The combination of these features is performed on the cost matrix level.
In common TbD based methods, the association relies on a cost matrix $C^{n\times m}$, where $n$ is the number of active trajectories (tracks) and $m$ is the number of detected objects. An element $C_{i,j}$ represents the cost of matching the $i$-th active track to the $j$-th detected object. The main objective of association is to find bipartite matches where the global cost is minimum (i.e., usually solved by the Hungarian algorithm [14]). In principle, $C_{i,j}$ is based on a distance $d$ from a stored/predicted cue of the $i$-th active track to a calculated/observed cue of the $j$-th detected object and is calculated in FaceQSORT as follows,
@@ 2 FaceQSORT
Figure 1: Overview FaceQSORT association.
In DeepSORT, a matching cascade is applied in which the matching is performed in sequence, starting with the tracks that were correctly matched in the previous frame and ending with the tracks that have not been matched for the longest time. The matching cascade reduces the search space and could be beneficial when long occlusions are allowed. An overview of the core association procedure of FaceQSORT is depicted in Figure 1. For all unmatched detections (after the matching cascade), an IoU-based association (as proposed in SORT [15]) is performed as a fallback routine. The general steps that are performed in FaceQSORT for each frame are summarized in Algorithm 1.
Algorithm 1 FaceQSORT at frame $t$
@@ 2.1 Computational Complexity
As shown in Algorithm 1, FaceQSORT essentially comprises the matching cascade, the IoU fallback matching, the feature updating and the position prediction. In general, the complexity of finding bipartite matches using the Hungarian algorithm is $O(|\Phi|^{2}\times|\Psi|)$. This is also the worst case with cascade matching. In the considered scenario, it is very likely that the detected face $\Psi$ in frame $t$ were also present in frame $t-1$. Thus, it is reasonable that the complexity can be reduced by the matching cascade (i.e., to $O((\nicefrac{{|\Phi|}}{{x}})^{2}\times\Psi)$). For all unmatched detections after the matching cascade, the complexity of the IoU fallback matching is in worst case again $O(|\Phi|^{2}\times|\Psi|)$. The complexity for updating the memory for biometric and appearance features ($\Phi_{bio}$ and $\Phi_{app}$) is $O(|\Psi|)$ in each case. Finally, predicting the positions in frame $t+1$ using the Kalman filter has basically a complexity of $O(\Psi)$. Thus, the overall complexity of FaceQSORT is,
@@ 4 Experimental Evaluation
FaceQSORT is evaluated based on the proposed PLUSFiaQ (see section 3) and the ChokePoint[47] dataset. In the ChokePoint dataset, people moving through a portal (chokepoint) are recorded. The dataset comprises sequences from 18 different runs (i.e., different portal, session or walking direction), with 3 different camera perspectives available for each run (i.e., $18\times 3=54$ available sequences). Of these 54 sequences, only 6 sequences show multiple persons per frame, which is similar to the assumed scenario in this paper. These 6 sequences are therefore used for evaluation. However, no annotations are available for these 6 sequences and are created accordingly. In the ChokePoint dataset, people are reflected in the open glass door as they walk through the portal. These refelections are annotated as non-target classes. The annotations created have the same format as described in section 3 and are made available along with the PLUSFiaQ dataset.
@@ 4.1 Tracker Implementation
FaceQSORT is implemented based on the StrongSORT implementation [^4] https://github.com/dyhBUPT/StrongSORT. The parameter $\alpha$ for the EMA feature updating strategy is set to $\alpha=0.9$. The weighting parameter $\beta$ (for weighting the contribution of biometric and appearance similarity compared to spatial distance) is set to 0.98. This gives the biometric and appearance similarity the most weight. $\theta_{pos}$ is set to 9.4877 (the 0.95 quantile of the chi-squared distribution with 4 degrees of freedom). These parameters are set according to [17]. The general cost threshold $\theta$ is set to 0.2 and the minimum IoU (fallback routine) to 0.3. For all experiments, $N_{init}=1$ and the maximum age parameter is set to $N_{max}=100$ frames (i.e., all tracks that were not successfully matched for more than 100 frames are deleted). With this setting, most occlusions are taken into account while avoiding interference from inter-track occlusions (as shown in Figure 5). As face detector yolov8555https://github.com/ultralytics/ultralytics with a confidence score threshold of 0.4 is used.
@@ 4.4 Experimental Results
In FaceQSORT, the selection of the weighting parameter $\lambda$ (combination of biometric and appearance features) and the association threshold $\theta$ is crucial. Thus, the parameter selection is evaluated in detail along with different face features and a different similarity metric (Euclidean distance instead of cosine similarity).
@@ 4.4.4 Matching Score Distribution
A multi-face tracker can be compared with a biometric system (e.g., a set of query images (detected faces) are matched against a database of enrolled identities (active tracks)). Matching is based on a similarity score, and in a perfect system, the genuine scores do not overlap with the impostor scores. In the context of FaceQSORT, the genuine scores represent the cosine similarity that result when a detected face is compared with the corresponding stored faces (active tracks). The imposter scores, on the other hand, represent the cosine similarities that are achieved when a detected face is compared with all other stored faces (tracks). Table 4 shows the mean and the standard deviation of the genuine and the imposter score distribution as well as their overlap (IoU). Considering the IoU, the most distinct face descriptor (lowest IoU) is SFace. Other face descriptors with a low IoU are Facenet512, VGG-Face, ArcFaceII and Dlib. These face descriptors (except Dlib) are also the face descriptors achieving the highest AssPr (see Figure 7). However, based on the IoU results, it is surprising that such low AssPr values are obtained with the Dlib face descriptor. A reason for the low AssPr values could be that, with a mean value of 0.07, the imposter scores are well below the threshold $\theta=0.2$ (applied to reject uncertain (imposter) matches). Thus, a closer look is taken at the selection of the threshold value $\theta$.
@@ 4.4.7 Compare Tracker
Table 8: Compare FaceQSORT with state-of-the-art multi-face tracker, where MC and IoU denote FaceQSORT wihtout matching cascade (MC) and/or IoU fallback matching.
@@
FaceQSORT MC 0.5940 0.7147 0.7025 0.0169 0.8274 0.8603 0.8700 0.0107 0.0405 0.1776 0.0875 0.0400
FaceQSORT IoU 0.5764 0.7033 0.6760 0.0187 0.6933 0.7863 0.7671 0.0283 0.0363 0.1676 0.0830 0.0494
FaceQSORT MC IoU 0.5589 0.6924 0.6666 0.0292 0.6697 0.7721 0.7521 0.0392 0.0359 0.1664 0.0812 0.0618
In table 8 FaceQSORT is compared with different state-of-the-art multi-face trackers. A comparison with StrongSORT including MC (i.e., when $\lambda$ is equal 1.0 or 0.0) is already performed in ‘1) Selection of a fixed $\lambda$’. The FaceQSORT results are obtained with SFace features, $\lambda=0.1$ and $\theta=0.2$. Implementations are available for the MCFDR [^7] https://github.com/yjwong1999/OpenVINO-Face-Tracking-using-YOLOv8-and-DeepSORT and UMA [^8] https://github.com/yinjunbo/UMA-MOT. However, no implementation is available for the OMTMCM multi-face tracker. For this reason, OMTMCM is carefully implemented according to the description in [42]. To verify the implementation, the results for the MusicVideo dataset [70] are also presented in table 8. The MusicVideo dataset consists of 8 music videos (i.e., Apink, BrunoMars, Darling, GirlsAloud, HelloBubble, PussycatDolls, Tara and Westlife) from YouTube, is publicly available [^9] https://sites.google.com/site/shunzhang876/eccv16-face-tracking and was used in [42] to evaluate the proposed OMTMCM multi-face tracker. Since music videos are unconstrained videos (with different scenes, moving camera, etc.), they do not correspond to the considered scenario in which people move towards a gate for which FaceQSORT is designed. For this dataset, OMTMCM performed best. However, in the considered sceneraio where people are moving towards a gate (PLUSFiaQ and ChokePoint), FaceQSORT clearly outperformed the other evaluated state-of-the-art trackers.
@@ 5 Ablation Study
The main components of FaceQSORT are the combination of biometric and appearance features, the matching cascade and IoU fallback matching. Running FaceQSORT with either biometric (i.e., $\lambda=1.0$) or appearance features (i.e., $\lambda=0.0$) has already been evaluated in ‘1) Selection of a fixed $\lambda$’ of subsection 4.4.
Table 9: Frames per second for FaceQSORT (FS), FaceQSORT without the matching cascade (FS MC) and OMTMCM [42].
@@
In the matching cascade (described in section 2), the tracks that were successfully matched in the previous frame are considered first. This can reduce the search space $\Phi$ and can be particularly beneficial if long occlusions are allowed (i.e., a large active track memory $\Phi$). For all experiments, occlusions up to 100 frames are allowed (i.e., $N_{max}=100$). In table 9, execution times in frames per second (fps) are reported for FaceQSORT with and without the matching cascade (MC) and compared with the re-implementation of a state-of-the-art tracker (i.e., OMTMCM [42]). All times were recorded on a standard desktop PC (i.e., Intel Core i5 gen9) and without taking face detection and feature extraction into account. It can be seen that the fps increases slightly when MC is applied. This is also reflected by the observation that on average less than half of all stored active tracks are required to successfully match all detected faces. However, applying the MC carries the risk that not the best matches are found. For example, it could happen that the best match (lowest cost) for a detected face is with a track that could not be matched in the last frames (e.g., a person returning from an occlusion), but the detected face is matched with a track that is tried in the MC before. The achieved tracking performance scores for FaceQSORT (with SFace features, $\lambda=0.1$ and $\theta=0.2$) without the matching cascade (MC) are reported in table 8. Compared to results when using the matching cascade (table 8 first row) a performance decrease can be observed.
IoU fallback matching is applied to all unmatched detections after the matching cascade, including matching with tentative tracks (only confirmed tracks are considered in the matching cascade). To achieve the reported tracking performance of FaceQSORT (table 8 first row), IoU matching is performed for 3.91%, 4.9% and 8.42% of all detected faces of the PLUSFiaQ, ChokePoint and MusicVideo dataset respectively. When disabling IoU fallback matching (IoU) the performance decreases significantly (see table 8). The performance achieved when FaceQORT is applied without matching cascade and IoU fallback matching is reported in table 8 (last row).